OpenStudy (anonymous):

x+3/x+2-x/x2-4=x/x-2

myininaya (myininaya):

multiply by x^2-4 giving you (x+3)(x-2)+(2-x)=x(x+2) assuming i read your question correctly

OpenStudy (anonymous):

no. x+3 divided by x+2-x divided by x2-4=x divided by x-2

OpenStudy (radar):

x=-3